OpenStudy (anonymous):

For a certain event, 812 tickets were sold, for a total of $1912. If students paid $2 per ticket and nonstudents paid $3 per ticket, how many student tickets were sold

OpenStudy (cwrw238):

let number student tickets = x and non students tickets = y x + y = 812 also 2x + 3y = 1912 solving these 2 equations gives your answer

OpenStudy (cwrw238):

can u do this?

OpenStudy (cwrw238):

multiply 1st equation by 2: 2x + 2y = 1624 2x + 3y = 1912 subtract first equation for second: y = 1912 - 1624 = 288 substitute for y in first equation x + 288 = 812 x = 812 - 288 = 524 so theres 524 students and 288 non student tickets sold.